OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope and containing the midpoint of the segment whose endpoints are (2, -3) and (-6, 5).

OpenStudy (zale101):

For the midpoint, we have to us this formula ((x1+x2)/2,(y1+y2)/2)

OpenStudy (anonymous):

Was the slope given? It loks like you forgot to add it or no?

OpenStudy (anonymous):

-4,6

OpenStudy (anonymous):

Wait where did you get 9 from? zale101

OpenStudy (zale101):

oh i'm sorry lol

OpenStudy (zale101):

So here we have ((2+(-6))/2,(-3+5)/2) =

OpenStudy (anonymous):

its ok i messed up too. it's 2, 3/2 or 7/2?

OpenStudy (zale101):

((2+(-6))/2= -2 (-3+5)/2= 1

OpenStudy (anonymous):

oh ok. what's next?

OpenStudy (zale101):

the slope is (change in y)/(change in x)

OpenStudy (zale101):

or (y2-y1)/(x2-x1)

OpenStudy (zale101):

So (5+(-3))/(-6+2) = ?

OpenStudy (anonymous):

averages?

OpenStudy (zale101):

Do you know how to plug in the numbers into the formula's, like what i did here for example (5+(-3))/(-6+2) = ?

OpenStudy (zale101):

solve this for me (5+(-3))/(-6+2) = ?

OpenStudy (anonymous):

sooo -3-5/2+6 = -8/8?

OpenStudy (zale101):

which equals -1

OpenStudy (anonymous):

so then after that i just do y-y1= -1(x1-x. which is y+3=-1(x-5)?

OpenStudy (zale101):

1 and -2 are the midpoints

OpenStudy (anonymous):

so (y-1)=-1(x=2)?

OpenStudy (zale101):

this means the midpoint is (-2,1)

OpenStudy (zale101):

x=-2 y=1

OpenStudy (anonymous):

so then what comes after finding the midpoint

OpenStudy (zale101):